OpenStudy (anonymous):

Find the three arithmetic means between -2 and 10.

OpenStudy (unklerhaukus):

add the terms together, then divide this quantity by the number of terms

OpenStudy (anonymous):

from -2 to 10 ?

OpenStudy (unklerhaukus):

what other data do you have?>

OpenStudy (anonymous):

Find the three arithmetic means between -2 and 10. Answer 2, 4, 7 1, 5, 7 1, 4, 7 2, 5, 8

OpenStudy (anonymous):

only tht

OpenStudy (unklerhaukus):

How can you get more than one arithmetic means from the same data?

OpenStudy (anonymous):

idk:(

hartnn (hartnn):

3 arithmetic means between a,b are x,y,z then a,x,y,z,b are in AP

hartnn (hartnn):

so lets first find the middle arithmetic mean 'y' a = -2 and b= 10 y = (a+b)/2

hartnn (hartnn):

now 'x' is the arithmetic mean between 'a' and 'y' x= (a+y)/2

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

no wait nevermind... idk

hartnn (hartnn):

z is the arithmetic mean between y and b z = (y+b)/2

hartnn (hartnn):

you mean y= 4 ?

OpenStudy (anonymous):

i think so

hartnn (hartnn):

thats correct :) find x and z too!

OpenStudy (anonymous):

wait how do i know whats x or y

hartnn (hartnn):

'x' is the arithmetic mean between 'a' and 'y' x= (a+y)/2 you already know y now :)

hartnn (hartnn):